# Learning When to Think: Shaping Adaptive Reasoning in R1-Style Models via Multi-Stage RL

### v2

@@ Abstract
Large reasoning models (LRMs) are proficient at generating explicit, step-by-step reasoning sequences before producing final answers. However, such detailed reasoning can introduce substantial computational overhead and latency, particularly for simple problems. To address this overthinking problem, we explore how to equip LRMs with adaptive thinking capabilities—enabling them to dynamically decide whether or not to engage in explicit reasoning based on problem complexity. Building on R1-style distilled models, we observe that inserting a simple ellipsis ("…") into the prompt can stochastically trigger either a thinking or no-thinking mode, revealing a latent controllability in the reasoning behavior. Leveraging this property, we propose AutoThink, a multi-stage reinforcement learning (RL) framework that progressively optimizes reasoning policies via stage-wise reward shaping. AutoThink learns to invoke explicit reasoning only when necessary, while defaulting to succinct responses for simpler tasks. Experiments on five mainstream mathematical benchmarks demonstrate that AutoThink achieves favorable accuracy–efficiency trade-offs compared to recent prompting and RL-based pruning methods. It can be seamlessly integrated into any R1-style model, including both distilled and further fine-tuned variants. Notably, AutoThink improves relative accuracy by 6.4% while reducing token usage by 52% on DeepSeek-R1-Distill-Qwen-1.5B, establishing a scalable and adaptive reasoning paradigm for LRMs.
@@ 2.2 Prompting Alone Does Not Enable Difficulty-Aware Thinking
A decreasing no-thinking rate along the difficulty axis reflects a desirable reasoning pattern, in which the model allocates effort based on task difficulty. However, this behavior does not emerge from prompting alone. Even with diverse prompt designs (Appendix A), the model failed to exhibit difficulty-aware reasoning. Yet prompt-only control suffers from a core limitation: without feedback, the model lacks a mechanism to learn when the thinking process is needed.
@@ 3.1 Stage 1: Preventing Mode Collapse by Batch Reward Balance
Let $z\in[0,1]$ denote the proportion of thinking trajectories in a training batch, and $1-z$ the no-thinking proportion. A target balance ratio $\gamma\in(0,1)$ and penalty slope $\lambda\geq 0$ control the strength of adjustment. For thinking and no-thinking samples, we compute soft penalty factors:
@@ 4.1 Setup
Table 1: Accuracy, Token Usage, and Efficiency Comparison Across Methods.
@@ 4.2 Main Results
[…]
@@ 4.4 In-Depth Behavioral and Efficiency Analysis
[…]
@@
We further conduct additional analyses on more base models, reasoning behaviors, training cost, and case studies. Details are presented in Appendix B due to space limitations.
@@ 5 Related Works
Reinforcement fine-tuning (RFT) has been widely adopted to improve the reasoning ability of LLMs [16, 11, 7, 9]. Recent work on RL for LLMs has focused on improving the efficiency and effectiveness of large-scale RL training. Key techniques decoupling the clipping mechanism and introducing dynamic group sampling [41], mitigating value bias over long sequences [43, 42], difficulty-aware advantage reweighting [45], and designing minimal-form credit assignment strategies for rewards [4]. In addition, RFT has been shown to explicitly promote self-verification and self-correction behaviors [25, 17], while also supporting optimization of test-time compute [22]. Multi-stage, context-length-extended RL further amplifies the long-chain reasoning ability of R1-style models [15, 27]. In our work, RL is applied to train R1-style models to adaptively control their reasoning behavior, enabling selective thinking guided by multi-stage reward shaping.
@@
While RFT improves performance, it may induce overthinking, causing models to generate overly verbose reasoning with limited benefit [28, 12]. [3] address overthinking in R1-style models by using self-generated short CoT as positive signals in DPO, encouraging concise reasoning. Recent studies have shown that inserting pseudo-thinking cues into R1-style prompts [18], or manually controlling reasoning based on problem difficulty [8, 35, 13], can suppress the model’s thinking behavior, but resulting in reduced performance. Other studies approach the problem from different perspectives: supervised fine-tuning (SFT) with short CoT responses [40, 19], incorporate response length–aware rewards in RFT [1, 39, 10, 6, 14], or leverage smaller models guide larger ones toward faster reasoning [13, 32]. Inspired by these findings, we first design a minimal prompt that elicits random thinking behavior, then apply multi-stage RL to guide the model to think adaptively based on task difficulty, without using external signals or teacher models.
@@ 6 Conclusion & Limitations
Appendix A Additional Prompt Variants
@@ B.1 Additional Results on Skywork-OR1-Math-7B
B.2 Reasoning Behavior Profiling
B.3 Training Cost Comparison
@@
B.4 Case Study

### v3

@@ Abstract
Large reasoning models (LRMs) are proficient at generating explicit, step-by-step reasoning sequences before producing final answers. However, such detailed reasoning can introduce substantial computational overhead and latency, particularly for simple problems. To address this overthinking problem, we explore how to equip LRMs with adaptive thinking capabilities, enabling them to dynamically decide whether to engage in explicit reasoning based on problem complexity. Building on R1-style distilled models, we observe that inserting a simple ellipsis ("…") into the prompt can stochastically trigger either a thinking or no-thinking mode, revealing a latent controllability in the reasoning behavior. Leveraging this property, we propose AutoThink, a multi-stage reinforcement learning (RL) framework that progressively optimizes reasoning policies via stage-wise reward shaping. AutoThink learns to invoke explicit reasoning only when necessary, while defaulting to succinct responses for simpler tasks. Experiments on five mainstream mathematical benchmarks demonstrate that AutoThink achieves favorable accuracy–efficiency trade-offs compared to recent prompting and RL-based pruning methods. It can be seamlessly integrated into any R1-style model, including both distilled and further fine-tuned variants. Notably, AutoThink improves relative accuracy by 6.4% while reducing token usage by 52% on DeepSeek-R1-Distill-Qwen-1.5B, establishing a scalable and adaptive reasoning paradigm for LRMs.
@@ 2.2 Prompting Alone Does Not Enable Difficulty-Aware Thinking
A decreasing no-thinking rate along the difficulty axis reflects a desirable reasoning pattern, in which the model allocates effort based on task difficulty. However, this behavior does not emerge from prompting alone. Even with diverse prompt designs (Appendix A.2), the model failed to exhibit difficulty-aware reasoning. Yet prompt-only control suffers from a core limitation: without feedback, the model lacks a mechanism to learn when the thinking process is needed.
@@ 3.1 Stage 1: Preventing Mode Collapse by Batch Reward Balance
Let $z\in[0,1]$ denote the proportion of thinking trajectories in a training batch, and 1z1-z the no-thinking proportion. A target balance ratio $\gamma\in(0,1)$ and penalty slope $\lambda\geq 0$ control the strength of adjustment. For thinking and no-thinking samples, we compute soft penalty factors:
@@ 4.1 Setup
Table 1: (Main Results) Accuracy, Token Usage, and Efficiency Comparison Across Methods.
@@ 4.2 Main Results
$\text{E-F1}=\left(2\cdot\frac{\Delta_{\text{acc}}\cdot\Delta_{\text{len}}}{\Delta_{\text{acc}}+\Delta_{\text{len}}}\right)\ \text{ if }\text{acc}>\text{acc}_{\text{std}}\text{ and }\text{len}<\text{len}_{\text{std}};\ \text{ else }0$
@@ 4.4 In-Depth Behavioral and Efficiency Analysis
We analyze how the trained model responds to the standard and forced-no-think prompts. The forced-no-think prompt is defined as $\n...\n</think>\n\n$, which builds upon the ellipsis prompt but enforces an immediate termination of the thinking phase. The results of Distill-1.5B-AutoThink are presented in Table 4.4. As expected, the standard prompt induces longer reasoning traces and achieves higher accuracy, while the forced no-think prompt reduces token usage at the cost of slight performance degradation. These findings suggest that AutoThink has learned to internally compress its reasoning when appropriate, while retaining the ability to conditionally invoke reasoning via prompting.
@@
Reasoning Behavior Profiling
Generality Beyond Mathematical Reasoning
To investigate whether AutoThink generalizes beyond mathematical reasoning, we additionally evaluate our models on three non-mathematical benchmarks: (i) GPQA for scientific multi-hop reasoning, (ii) MMLU for general multi-task language understanding, and (iii) Live-Code-Bench for code generation (20250727 release).
As shown in Table 4.4, AutoThink retains competitive accuracy while reducing token usage, indicating that adaptive reasoning behaviors extend beyond math tasks. Stage 2 even surpasses the baseline in accuracy while halving response length, highlighting the transferability of our approach to diverse domains.
Table 4: Performance of AutoThink on non-math benchmarks. Each cell shows Accuracy (%) / Avg. Length.
GPQA MMLU Live-Code-Bench Avg
Distill-1.5B 35.1 / 10026 49.5 / 2727 25.2 / 13372 36.6 / 8708
AutoThink-Stage1 31.5 / 8889 47.7 / 1190 23.8 / 5653 34.3 / 5244
AutoThink-Stage2 37.1 / 8617 48.8 / 1743 24.2 / 9647 36.7 / 6669
AutoThink-Stage3 35.7 / 5659 48.8 / 1300 24.9 / 9054 36.5 / 5337
@@
We further conduct additional analyses on more base models, hyperparameters, training cost, and case studies. Details are presented in Appendix B due to space limitations.
@@ 5 Related Works
Reinforcement fine-tuning (RFT) has been widely adopted to improve the reasoning ability of LLMs [32, 13, 9, 11, 7, 34]. Recent work on RL for LLMs has focused on improving the efficiency and effectiveness of large-scale RL training. Key techniques decoupling the clipping mechanism and introducing dynamic group sampling [44], mitigating value bias over long sequences [46, 45], difficulty-aware advantage reweighting [48], model ensembling [8] and designing minimal-form credit assignment strategies for rewards [4]. In addition, RFT has been shown to explicitly promote self-verification and self-correction behaviors [26, 18], while also supporting optimization of test-time compute [23]. Multi-stage, context-length-extended RL further amplifies the long-chain reasoning ability of R1-style models [17, 28]. In our work, RL is applied to train R1-style models to adaptively control their reasoning behavior, enabling selective thinking guided by multi-stage reward shaping.
@@
While RFT improves performance, it may induce overthinking, causing models to generate overly verbose reasoning with limited benefit [29, 14]. [3] address overthinking in R1-style models by using self-generated short CoT as positive signals in DPO, encouraging concise reasoning. [49] mitigate overthinking by training models to terminate with “I don’t know” on unsolvable problems. Recent studies have shown that inserting pseudo-thinking cues into R1-style prompts [19], or manually controlling reasoning based on problem difficulty [10, 38, 15], can suppress the model’s thinking behavior, but resulting in reduced performance. Other studies approach the problem from different perspectives: supervised fine-tuning (SFT) with short CoT responses [43, 20], incorporate response length–aware rewards in RFT [1, 42, 12, 6, 16], or leverage smaller models guide larger ones toward faster reasoning [15, 35]. Inspired by these findings, we first design a minimal prompt that elicits random thinking behavior, then apply multi-stage RL to guide the model to think adaptively based on task difficulty, without using external signals or teacher models.
@@ 6 Conclusion & Limitations
Acknowledgments and Disclosure of Funding
This work is supported by the Strategic Priority Research Program of Chinese Academy of Sciences under Grant XDA0480303, Young Scientists Fund of The State Key Laboratory of Multimodal Artificial Intelligence Systems ES2P100112, National Natural Science Foundation of China 62402252 and 62536003.
Appendix A Additional Definition and Prompts
A.1 Definition of No-Thinking
In R1-style models (e.g., DeepSeek-R1), Thinking refers to generating explicit, step-by-step reasoning traces enclosed within <think> $\cdots$ </think>, enabling reflection and backtracking. By contrast, we define the No-Thinking mode [19] as immediately closing the <think> tag without producing any substantive reasoning, e.g., <think> </think> before moving to the final answer. This phenomenon, also referred to as Non-Thinking in the Qwen3 Technical Report [30], often emerges under our ellipsis prompt, which stochastically toggles the model between reasoning and shortcut modes and thus serves as a lightweight control signal for studying adaptive reasoning behaviors.
A.2 Additional Prompt Variants
@@ B.1 Additional Results on Skywork-OR1-Math-7B
B.2 Additional Prompt Evaluation on Qwen3
We further extend our study to the Qwen3-8B [30] model by applying the proposed ellipsis prompting and adaptive training strategy. Table 7 reports accuracy and average token length across benchmarks. The results show that ellipsis prompting encourages a non-negligible amount of no-thinking behavior; however, this tendency does not align perfectly with task difficulty (e.g., AIME problems are significantly harder than MATH500, yet elicit a lower thinking rate). Together with observations on Skywork-OR1-Math-7B (where ellipsis prompting induced only $\sim$0.5% no-thinking behavior), these findings suggest that the AutoThink strategy can also induce autonomous reasoning behavior in Qwen3, with $\sim$13% occurrence of no-thinking responses under ellipsis prompting.
Table 7: Results of Qwen3-8B with different prompts. Each cell shows Accuracy (%) / Avg. Length.
Qwen3-8B MATH500 Minerva Olympiad AIME24 AMC23 Avg
Standard Prompt 97.0 / 5351 53.3 / 7010 73.5 / 11342 86.7 / 14690 88.1 / 10343 79.7 / 9747
Ellipsis Prompt 96.4 / 5109 49.5 / 5315 70.9 / 9891 68.3 / 13349 88.9 / 9858 74.8 / 8704
No-Thinking Prompt 84.1 / 1104 41.2 / 639 50.8 / 2860 26.3 / 6518 60.1 / 2913 52.5 / 2807
Ellipsis Prompt: Thinking Rate 96.9% 67.5% 89.0% 87.5% 96.2% 87.4%
B.3 Hyperparameter Sensitivity
The three-stage framework is intentionally designed to be modular and interpretable, with each stage serving a distinct and simple role: (i) Stage 1 introduces a batch-wise reward balance to prevent mode collapse between thinking and no-thinking behaviors; (ii) Stage 2 focuses purely on reinforcing accuracy within each mode without additional reward shaping; (iii) Stage 3 adds length-aware shaping to encourage brevity for correct responses and elaboration for incorrect ones.
Among these stages, only Stage 1 and Stage 3 involve reward shaping beyond naive correctness. Even in these cases, the formulations remain straightforward and principled. Specifically, Stage 1 balances the modal ratio using a linear penalty controlled by hyperparameters $\gamma$ and $\lambda$, which are set to simple default values rather than finely tuned. As illustrated in Figure 3, the resulting reward curve naturally exhibits a symmetric form. Stage 3 reuses shaping terms ($\alpha$, $\beta$) inspired by GRPO-LEAD, again without introducing any ad hoc modifications.
To further examine the robustness of these choices, we conduct sensitivity analyses of $\gamma$, $\lambda$, and $\alpha$ during training, with results summarized below.
Stage 1 Parameters ($\gamma$, $\lambda$).
In Stage 1, we balance the modal ratio between thinking and no-thinking behaviors using a linear penalty controlled by $\gamma$ and $\lambda$. These hyperparameters were not carefully tuned but set to commonly used values. To illustrate their effect, Table 8 reports the average thinking rate at steps 100 and 200 during training. Increasing $\gamma$ encourages more thinking trajectories, while larger $\lambda$ enforces stricter adherence to the target balance.
Table 8: Thinking rate (%) at checkpoints under different values of $\gamma$ and $\lambda$.
$\gamma{=}0.5,\lambda{=}2$ $\gamma{=}0.2,\lambda{=}2$ $\gamma{=}0.8,\lambda{=}2$ $\gamma{=}0.5,\lambda{=}1$ $\gamma{=}0.5,\lambda{=}4$
Thinking-Rate@step100 62.4 57.8 99.9 71.4 51.7
Thinking-Rate@step200 54.2 51.4 100.0 61.2 48.3
Stage 3 Parameters ($\alpha$, $\beta$).
In Stage 3, the shaping terms ($\alpha$, $\beta$) control the rate of reward decay/growth with respect to response length. Table 9 shows response length at checkpoints under different $\alpha$ values with $\beta$ fixed at 0.05. Larger $\alpha$ accelerates length decay for correct responses, while smaller $\alpha$ relaxes the penalty.
Table 9: Response length under different $\alpha$ values ($\beta=0.05$).
$\alpha=0.05$ $\alpha=0$ $\alpha=0.1$
Response-Length@step100 4734 6174 3623
Response-Length@step200 4120 6322 2894
Discussion.
These results confirm that the shaping functions behave as intended: $\gamma$$\lambda$ modulate the balance between modes in Stage 1, and $\alpha$$\beta$ regulate the brevity of responses in Stage 3. Importantly, the overall training trends remain consistent with the main results, demonstrating the robustness of AutoThink without extensive hyperparameter tuning. Thus, while the overall pipeline appears multi-stage, each stage was deliberately designed with minimal tuning and clear interpretability. Looking ahead, it may be possible to unify these stages through a more holistic reward formulation, enabling the model to learn adaptive reasoning behavior within a single-stage process. We leave this as a promising direction for future work.
B.4 Training Cost Comparison
@@
Table B.4 provide estimated GPU-hour costs for all methods using Distill-R1-1.5B as the base model, AutoThink operates within the same order of compute as concise baselines such as ThinkPrune and ShorterBetter, yet achieves notably stronger performance. In contrast, DeepScaleR, which primarily aims to maximize performance, requires more than $3\times$ higher compute due to its longer context length and increased RL iterations.
Table 11: Estimated GPU-hour cost of training Distill-R1-1.5B with different methods on H100.
Method GPU Hours Avg ACC / Length
ThinkPrune-iter-2K $\sim$400 49.2 / 3368
ShorterBetter $\sim$200 44.7 / 1915
Distill-1.5B-AutoThink $\sim$700 51.7 / 5108
DeepScaleR $\sim$2200 56.7 / 5817
DeepScaleR-1.5B-AutoThink $\sim$250 57.3 / 5277
Appendix C Addressing Potential Challenges
While AutoThink demonstrates robust improvements, several open challenges remain in the area of reasoning control. We outline possible directions for addressing these limitations below:
Token-Budget Control.
Token-budget constraints have been partially explored in prior works [1, 42], where budget-aware reward functions penalize excessively long completions. Such formulations can be readily integrated with AutoThink in a plug-and-play manner to enforce global compute budgets and further improve efficiency.
Dataset Noise.
The presence of noise in large-scale reasoning datasets can hinder training efficiency. Prior studies [28] suggest that curriculum learning or filtering samples by correctness or difficulty can improve learning quality. These strategies are orthogonal to our reward design and could be combined with AutoThink to further enhance robustness.
Reward Hacking.
A common issue is reward hacking, where the model continues reasoning after the </think> tag. This can be mitigated by explicitly penalizing reasoning-related patterns outside the <think> span, or by rewarding clean separation between thought and final answer. Both strategies can be incorporated into future iterations of our reward function.
Overall, we view these solutions as complementary and composable with our framework. Future work will explore tighter integration of these mechanisms to provide a more comprehensive solution to reasoning control.
Appendix D Case Study